# Text Generation Beyond Discrete Token Sampling

### v2

@@ 3.2 Bayesian Input Construction with MoI
To capture the distribution information, a naive idea might simply be to directly mix the inputs according to the output distribution, setting $w_{t,i}=p_{t,i}$. However, this approach only treats the token distribution as the input and neglects the sampled next token. In Section 6.1, we experiment with this approach (referred to as Direct Mixture) and find that it leads to performance degradation in most cases.
@@ 4.2 Estimating Mixing Weight
Let $\boldsymbol{p}_{t}\in\Delta^{V-1}$ be the next-token distribution at step $t$ and let $y_{t,i}\in\{0,1\}$ indicate the sampled token ($y_{t,i}=1$ iff token $i$ is chosen). We estimate the mixing weights $\boldsymbol{w}_{t}$ by Bayesian posterior inference in a Dirichlet–Multinomial model.
@@
The sampled token contributes a single pseudo-count whose weight increases as confidence rises:
@@
The hyperparameter $\beta$ controls the concentration over mixing weight. Smaller values emphasize more on output distributions, while larger values highlight more the sampled output token. The effect of $\beta$ is easier to observe in Eq. 5. We also conduct an analysis of $\beta$’s empirical effect in Section 7.1.
@@
Eq. 5 smoothly interpolates between the distribution ($w_{t,i}\rightarrow p_{t,i}$ when $H\rightarrow 1$) and the one-hot token ($w_{t,i}\rightarrow y_{t,i}$ when $H\rightarrow 0$), thereby reconciling distributional and discrete evidence in a single principled estimator.
The complete procedure for computing the mixture of inputs is summarized in Algorithm 1.
@@ 5 Experimental Setup
0:  Sampling distribution $\boldsymbol{p}_{t}$, sampled token $\boldsymbol{y}_{t}$, hyperparameter $\beta$, and embeddings $\{e_{i}\}_{i=1}^{V}$.   1. Compute entropy $H$ with Eq. 2   2. Compute mixing weight $w_{t,i}$ with Eq. 5   return  $\boldsymbol{h}_{t}=\sum_{i}w_{t,i}\,\boldsymbol{e}_{i}$
@@ 5.2 Models
Gemma-3-27B [33] is part of Google’s Gemma 3 family—multimodal (text + image) models with 128 K token context windows and an integrated SigLIP vision encoder. The 27B variant is instruction-tuned for chat and reasoning.
DAPO-Qwen-32B [34] is a customized version of Qwen2.5-32B [35] that incorporates Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO), which stabilizes and scales RL for long chain-of-thought reasoning. This model is designed to encourage faithful and step-consistent reasoning trajectories.
@@ 5.4 Hyperparameter Settings
[…]
@@ 6.2 Hyperparameter Importance Analysis
To understand which factors most strongly influence reasoning performance, we analyze three key hyperparameters: $\beta$, top-p, and temperature. This analysis spans four LLMs and two mathematical reasoning tasks, with multiple runs and grid search over the hyperparameter space, as described in Section 5.4.
Fig. 2 provides two complementary perspectives on hyperparameter importance. The left plot tracks expected performance gain when optimizing each parameter individually through best-of-N-shots tuning, with experiment setup explained in Section C.1. As N increases from 1 to 15, $\beta$ consistently yields the highest gains, reaching nearly 7.5% improvement at N=15, while top-p and temperature plateau at approximately 6.0-6.5%. This separation becomes particularly pronounced after N=10, suggesting that $\beta$’s impact grows with more extensive search.
The right panel quantifies each parameter’s importance through random forest regression analysis, with experiment setup explained in Section C.2. With inputs as hyperparameters and accuracy as the target, this reveals $\beta$ as the dominant factor (importance score of 0.41), followed by top-p (0.32) and temperature (0.27).
@@ 7.1 Task-Dependent Optimal Mixing Strategies
Different reasoning tasks may benefit from varied degrees of distribution mixing. To investigate this phenomenon, we analyze the parameter sensitivity of two distinct benchmark types: AIME (requiring advanced mathematical reasoning) and Count Down 4 (demanding extensive combinatorial enumeration). Fig. 3 visualizes how performance varies with the mixing parameter $\beta$ across four LLMs, showing the deviation from each task’s global mean accuracy.
@@ 7.2 Case Study: Linear Prompt Blending with Various Lengths
Although our main experiments focus on token-by-token blending at generation time, we also investigate whether a similar blending strategy applied to instruction prompts of varying lengths can boost performance. To this end, we perform 10-shot in-context learning on five sentiment analysis benchmarks using three medium-sized LLMs, building on prior work showing that prompt wording and structure have a major impact on classification accuracy [15].
We assembled three prompt pools: (1) binary sentiment analysis on Rotten Tomatoes [36], SST2 [37], and IMDB [38], consisting of 96 prompts of length 3–16 words (mean 7.57); (2) 6-class emotion classification on the Emotion dataset [39], with 32 prompts of length 3–15 words (mean 7.27); and (3) 3-class financial sentiment on the Financial Phrasebank [40], comprising 40 prompts of length 2–14 words (mean 6.51).
Our blending procedure first linearly extrapolates each prompt’s embedding to the maximum length in its pool and then averages these fixed-length embeddings to form a single “blended” prompt representation. This approach integrates semantic nuances from all constituent prompts while preserving their instructional intent.
Table 2 reports 10-shot accuracy under the expectation over randomly drawn single-prompt, the blended-prompt, and the absolute gain over baseline. Across most benchmarks and models, linear prompt blending consistently outperforms random single-prompt selection, further demonstrating that embedding-space mixing can be highly effective for boosting LLMs’ capacity.
@@ 7.3 Throughput Analysis
The mixing weight calculation is lightweight and efficient. We perform a throughput analysis, shown in Table 3, to examine the runtime overhead added by MoI. We measure generation statistics for solving the Count Down 4 task, and we record the average input and output throughput over 5 runs.
@@ Appendix A Comparing over Hyperparameter Grid Search
Best-case: each method is run with its single best-performing hyperparameter configuration, Fig. 4 summarizes the results.
Grid-average: performance is averaged across all combinations in the hyperparameter grid (see details in Section 5.4). Fig. 5 provides these averages and confirms that the gains are not an artifact of cherry-picking one lucky setting.
Figure 4: We show a comparison of distributions of evaluation results across the best top-$p$ and temperature hyperparameter for baseline and with MoI. The results indicate strong performance gain brought by incorporating the sampling distribution in the generation process.
Figure 5: We show a comparison of distributions of evaluation results across all top-$p$ and temperature hyperparameters. The results indicate almost universal performance gain across average hyperparameter settings.
@@ Appendix B Generalization of a single hyper-parameter setting across tasks
Table 4: Additional results on four benchmarks with four large language models. For every (method, model) we tune the decoding hyperparameters on AIME and CountDown 4, fix the best setting, and reuse it when assessing GPQA-Diamond and LiveCodeBench. Accuracy (%) is reported on AIME, Count Down 4, GPQA-Diamond, and pass@1 is used on LiveCodeBench. Standard uses conventional sampling, that is temperature–scaled nucleus sampling and MoI is our full approach.
@@
A practical concern when adding new decoding hyperparameters is whether the values tuned on one task will transfer to others. To investigate this, we conduct a grid search over the hyperparameters described in Section 5.4. For each (method, model) pair, we keep the configuration that maximizes the average accuracy on AIME & CountDown 4 and freeze it for the remainder of the study.
Table 4 shows the results on the held-out GPQA-Diamond and LiveCodeBench. The single-tuned setting consistently outperforms Standard nucleus sampling on 15 of the 16 (model, task) pairs.
@@
Appendix C Additional Setups
C.1 Best-of-N Analysis Setup
@@
To measure how quickly a limited tuning budget yields performance gains, we simulate a best-of-$N$ random search for $N\!=\!1,\dots,15$. For every model-task pair in AIME and Count Down 4 we start from the complete Cartesian grid of hyperparameters in Table 4. At each Monte-Carlo replicate we uniformly draw $N$ distinct values for a single target hyperparameter ($\beta$, top-p, or temperature) while keeping the other two at their default settings, retrieve the corresponding validation accuracies that were pre-computed during the main grid search, and record the improvement of the best sampled configuration over the initial draw. Repeating this procedure 256 times and averaging across the four LLMs produces the curves in Fig. 2.
C.2 Random-Forest Regression Analysis Setup
@@
Every completed grid-search run — defined by a specific model, task, and random seed—serves as one training example for a model-agnostic importance analysis. Each example is encoded by the triple $(\beta,\text{top-}p,T)$ and labeled with its accuracy. We fit a RandomForestRegressor from scikit-learn [43] with 100 trees that have unrestricted depth. Impurity-based Gini importances rank the hyperparameters as $\beta$ (0.41), top-p (0.32), and temperature (0.27), as shown in Fig. 2.
@@
Table 5: Hyperparameter configuration by task. AIME and Count Down 4 use grid-search ranges; GPQA-Diamond and LiveCodeBench share a single universal setting.
@@
$\beta$$\{\,\tfrac{1}{4},\tfrac{1}{2},1,2,4,8\}$$\{\,\tfrac{1}{4},\tfrac{1}{2},1,2,4,8\}$ $1$ $1$
@@
C.3 Setup for Case Study
@@
During inference, we prepend this blended vector to each of the 10-shot demonstrations and feed the embeddings directly to the model. For the single-prompt baseline we calculate the average accuracy of the prompts as the expectation of randomly choosing any single prompt, whereas the blended prompt is evaluated once; consequently, Table 2 reflects an identical number of forward passes per model. Because all other factors remain fixed, any performance difference isolates the effect of the prompt representation itself.
Appendix D Hyperparameters
Table 5 lists the full search space for AIME and Count Down 4, along with the universal settings used for GPQA-Diamond and LiveCodeBench. All searches use five random seeds; reported results are seed-averaged.
Appendix E Implementation Details

### v3

# Mixture of Inputs: Text Generation Beyond Discrete Token Sampling

@@ Abstract
footnotetext: Code is available at: https://github.com/EvanZhuang/mixinputs.
@@ 3.2 Bayesian Input Construction with MoI
To capture the distribution information, a naive idea might simply be to directly mix the inputs according to the output distribution, setting $w_{t,i}=p_{t,i}$. However, this approach only treats the token distribution as the input and neglects the sampled next token. In Section˜6.1, we experiment with this approach (referred to as Direct Mixture) and find that it leads to performance degradation in most cases.
@@ 4.2 Estimating Mixing Weight
Let $\boldsymbol{p}_{t}\in\Delta^{V-1}$ be the next‑token distribution at step $t$ and let $y_{t,i}\in\{0,1\}$ indicate the sampled token ($y_{t,i}=1$ iff token $i$ is chosen). We estimate the mixing weights $\boldsymbol{w}_{t}$ by Bayesian posterior inference in a Dirichlet–Multinomial model.
@@
The sampled token contributes a single pseudo‑count whose weight increases as confidence rises:
@@
The hyperparameter $\beta$ controls the concentration over mixing weight. Smaller values emphasize more on output distributions, while larger values highlight more the sampled output token. The effect of $\beta$ is easier to observe in Eq.˜5. We also conduct an analysis of $\beta$’s empirical effect in Section˜7.1.
@@
Eq.˜5 smoothly interpolates between the distribution ($w_{t,i}\rightarrow p_{t,i}$ when $H\rightarrow 1$) and the one‑hot token ($w_{t,i}\rightarrow y_{t,i}$ when $H\rightarrow 0$), thereby reconciling distributional and discrete evidence in a single principled estimator.
The complete procedure for computing the mixture of inputs is summarized in Algorithm˜1.
@@ 5 Experimental Setup
0:  Sampling distribution $\boldsymbol{p}_{t}$, sampled token $\boldsymbol{y}_{t}$, hyperparameter $\beta$, and embeddings $\{e_{i}\}_{i=1}^{V}$.
  1. Compute entropy $H$ with Eq. 2
  2. Compute mixing weight $w_{t,i}$ with Eq. 5
  return $\boldsymbol{h}_{t}=\sum_{i}w_{t,i}\,\boldsymbol{e}_{i}$
@@ 5.2 Models
Gemma-3-27B [33] is part of Google’s Gemma 3 family—multimodal (text + image) models with 128 K token context windows and an integrated SigLIP vision encoder. The 27B variant is instruction‐tuned for chat and reasoning.
DAPO-Qwen-32B [34] is a customized version of Qwen2.5-32B [35] that incorporates Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO), which stabilizes and scales RL for long chain‐of‐thought reasoning. This model is designed to encourage faithful and step-consistent reasoning trajectories.
@@ 5.4 Hyperparameter Settings
We perform 5 runs for all experiments and report the average. For AIME and Count Down 4, we perform hyperparameter grid search on baselines, Direct Mixture and MoI with $\beta\in\{\tfrac{1}{4},\tfrac{1}{2},1,2,4,8\}$, $T\in\{0.6,0.8,1\}$ and $\text{top-p}\in\{0.4,0.6,0.8,0.95\}$. We report the mean result of the best configuration for all three methods. We investigate the importance of these hyperparameters in Section˜6.2. For GPQA-Diamond and LiveCodeBench, we use the universal hyperparameter for all of them with $T=0.6,\text{top-p}=0.95,\beta=1$; more details can be found in Appendix˜F.
@@ 6.2 Hyperparameter Importance Analysis
To understand which factors most strongly influence reasoning performance, we analyze three key hyperparameters: $\beta$, top-p, and temperature. This analysis spans four LLMs and two mathematical reasoning tasks, with multiple runs and grid search over the hyperparameter space, as described in Section˜5.4.
Fig.˜2 provides two complementary perspectives on hyperparameter importance. The left plot tracks expected performance gain when optimizing each parameter individually through best-of-N-shots tuning, with experiment setup explained in Section˜E.1. As N increases from 1 to 15, $\beta$ consistently yields the highest gains, reaching nearly 7.5% improvement at N=15, while top-p and temperature plateau at approximately 6.0-6.5%. This separation becomes particularly pronounced after N=10, suggesting that $\beta$’s impact grows with more extensive search.
The right panel quantifies each parameter’s importance through random forest regression analysis, with experiment setup explained in Section˜E.2. With inputs as hyperparameters and accuracy as the target, this reveals $\beta$ as the dominant factor (importance score of 0.41), followed by top-p (0.32) and temperature (0.27).
@@ 7.1 Task-Dependent Optimal Mixing Strategies
Different reasoning tasks may benefit from varied degrees of distribution mixing. To investigate this phenomenon, we analyze the parameter sensitivity of two distinct benchmark types: AIME (requiring advanced mathematical reasoning) and Count Down 4 (demanding extensive combinatorial enumeration). Section˜7.3 visualizes how performance varies with the mixing parameter $\beta$ across four LLMs, showing the deviation from each task’s global mean accuracy.
@@ 7.2 Case Study: Linear Prompt Blending with Various Lengths
Although our main experiments focus on token‐by‐token blending at generation time, we also investigate whether a similar blending strategy applied to instruction prompts of varying lengths can boost performance. To this end, we perform 10‐shot in‐context learning on five sentiment analysis benchmarks using three medium-sized LLMs, building on prior work showing that prompt wording and structure have a major impact on classification accuracy [15].
We assembled three prompt pools: (1) binary sentiment analysis on Rotten Tomatoes [36], SST2 [37], and IMDB [38], consisting of 96 prompts of length 3–16 words (mean 7.57); (2) 6‐class emotion classification on the Emotion dataset [39], with 32 prompts of length 3–15 words (mean 7.27); and (3) 3‐class financial sentiment on the Financial Phrasebank [40], comprising 40 prompts of length 2–14 words (mean 6.51).
Our blending procedure first linearly extrapolates each prompt’s embedding to the maximum length in its pool and then averages these fixed‐length embeddings to form a single “blended” prompt representation. This approach integrates semantic nuances from all constituent prompts while preserving their instructional intent.
Table˜2 reports 10‐shot accuracy under the expectation over randomly drawn single‐prompt, the blended‐prompt, and the absolute gain over baseline. Across most benchmarks and models, linear prompt blending consistently outperforms random single‐prompt selection, further demonstrating that embedding‐space mixing can be highly effective for boosting LLMs’ capacity.
@@ 7.3 Throughput Analysis
The mixing weight calculation is lightweight and efficient. We perform a throughput analysis, shown in Section˜7.3, to examine the runtime overhead added by MoI. We measure generation statistics for solving the Count Down 4 task, and we record the average input and output throughput over 5 runs. To better compare the throughput, we picked the benchmarks where the generation length is about the same. The median difference between MoI-generated text and baseline-generated text is 1.7%.
@@ Conclusion
Acknowledgement
Our work is sponsored in part by NSF CAREER Award 2239440, NSF Proto-OKN Award 2333790, Sponsored Research Projects from companies like Cisco and eBay, as well as generous gifts from Google, Adobe, and Teradata. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and should not be interpreted as necessarily representing the views, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for government purposes not withstanding any copyright annotation hereon.
@@ Appendix A Comparing over Hyperparameter Grid Search
Best-case: each method is run with its single best-performing hyperparameter configuration, Fig.˜A1 summarizes the results.
Grid-average: performance is averaged across all combinations in the hyperparameter grid (see details in Section˜5.4). Fig.˜A2 provides these averages and confirms that the gains are not an artifact of cherry-picking one lucky setting.
Figure A1: We show a comparison of distributions of evaluation results across the best top-$p$ and temperature hyperparameter for baseline and with MoI. The results indicate strong performance gain brought by incorporating the sampling distribution in the generation process.
Figure A2: We show a comparison of distributions of evaluation results across all top-$p$ and temperature hyperparameters. The results indicate almost universal performance gain across average hyperparameter settings.
@@ Appendix B Generalization of a single hyper-parameter setting across tasks
Table A1: Additional results on four benchmarks with four large language models. For every (method, model) we tune the decoding hyperparameters on AIME and CountDown 4, fix the best setting, and reuse it when assessing GPQA-Diamond and LiveCodeBench. Accuracy (%) is reported on AIME, Count Down 4, GPQA-Diamond, and pass@1 is used on LiveCodeBench. Standard uses conventional sampling, that is temperature–scaled nucleus sampling and MoI is our full approach.
@@
A practical concern when adding new decoding hyperparameters is whether the values tuned on one task will transfer to others. To investigate this, we conduct a grid search over the hyperparameters described in Section˜5.4. For each (method, model) pair, we keep the configuration that maximizes the average accuracy on AIME & CountDown 4 and freeze it for the remainder of the study.
Table˜A1 shows the results on the held-out GPQA-Diamond and LiveCodeBench. The single-tuned setting consistently outperforms Standard nucleus sampling on 15 of the 16 (model, task) pairs.
@@
Appendix C Results on Instruction-following Tasks
We conducted additional experiments on MT-Bench [43], using the four larger models, and the two smaller models, Llama-3.1-8B [32] and Qwen-2.5-14B [35]. Below, we report the average performance for standard decoding (temperature=0.6, top-p=0.95) and our Mixture of Inputs (MoI) decoding ($\beta$=1,2,4, same decoding configuration), together with the relative improvements brought by MoI ($\Delta$). As summarized in Table˜A2, MoI consistently outperforms the baseline for all models.
We also noticed MoI reduced repetition in the response. Particularly, we measured the 3-gram and 4-gram repetition rate in the response for MT-bench and found MoI reduces the repetition on 72.9% of responses, reducing the average 4-grams repetition rate from 4.27% to 3.13%. We also provide a side-by-side comparison on a coding task as below (QwQ-32B).
With standard sampling (repetitive around HCA definition):
To address the problem of finding the highest common ancestor (HCA) of two nodes in a binary tree, we need to clarify the definition. Based on common terminology, the lowest common ancestor (LCA) is the deepest node shared by both paths from the nodes to the root. However, since the problem specifies "not LCA," we assume the HCA refers to the shallowest common ancestor (closest to the root) [content continues…]
With mixture of inputs sampling (reduced repetitiveness):
To find the highest common ancestor (HCA) of two nodes in a binary tree, we can utilize a recursive approach similar to finding the Lowest Common Ancestor (LCA). The HCA is the deepest node that is an ancestor of both given nodes. Here’s how to implement this: [content continues…]
Table A2: Comparison of conventional sampling (Baseline) and MoI scores across six large language models. For each model, we report the baseline score, MoI score, and the relative improvement ($\Delta$%) of MoI over the baseline.
Model Baseline Score MoI Score $\Delta$% (MoI vs Baseline)
QwQ-32B 9.25 9.51 +2.81%
Nemotron-Super-49B 9.41 9.48 +0.74%
Gemma-3-27B 9.05 9.38 +3.65%
DAPO-Qwen-32B 8.96 9.46 +5.54%
Llama-3.1-8B 8.24 8.65 +4.98%
Qwen-2.5-14B 8.87 9.32 +5.07%
Appendix D Statistical Robustness of the Results
We conducted formal significance testing. Specifically, we applied McNemar’s test to compare MoI against the baseline across all four benchmarks for each model. Table˜A3 reports the resulting p-values. These results validate that the observed gains are statistically significant in most cases.
To further address concerns about variance, we conducted an extensive 64-run evaluation on the AIME dataset across four models with the same configuration as in experiments of Table˜1. The results in Table˜A4 confirm consistent improvements from MoI.
Table A3: McNemar’s test $p$-values comparing model performance across four benchmarks. Lower values indicate statistically significant differences between baseline and MoI.
Model CountDown4 AIME GPQA-Diamond LiveCodeBench
QwQ-32B <0.001<0.001 <0.001<0.001 $0.003$ $0.04$
Nemotron-Super-49B <0.001<0.001 <0.001<0.001 <0.001<0.001 $0.02$
Gemma-3-27B <0.001<0.001 <0.001<0.001 $0.04$ <0.001<0.001
DAPO-Qwen-32B <0.001<0.001 $0.8$ $0.02$ <0.001<0.001
Table A4: Comparison of baseline and MoI performance across four large language models over 64 runs on AIME. Each result reports the mean ($\mu$), standard deviation ($\sigma$) and min-max score range. We observe consistent performance gain brought by MoI.
Model Baseline ($\mu$ $\pm$ $\sigma$) Baseline Range MoI ($\mu$ $\pm$ $\sigma$) MoI Range Gain
QwQ-32B 76.02 $\pm$ 2.51 72.22–82.22 77.66 $\pm$ 2.06 73.33–82.22 +1.64
Nemotron-Super-49B 54.80 $\pm$ 2.93 48.89–61.11 55.03 $\pm$ 3.54 44.44–62.22 +0.23
Gemma-3-27B 21.80 $\pm$ 2.34 16.67–27.78 24.91 $\pm$ 2.30 20.00–30.00 +3.11
DAPO-Qwen-32B 61.16 $\pm$ 2.50 56.67–66.67 62.57 $\pm$ 2.49 57.78–68.89 +1.41
Appendix E Additional Setups
E.1 Best-of-N Analysis Setup
@@
To measure how quickly a limited tuning budget yields performance gains, we simulate a best-of-$N$ random search for $N\!=\!1,\dots,15$. For every model-task pair in AIME and Count Down 4 we start from the complete Cartesian grid of hyperparameters in Table 4. At each Monte-Carlo replicate we uniformly draw $N$ distinct values for a single target hyperparameter ($\beta$, top-p, or temperature) while keeping the other two at their default settings, retrieve the corresponding validation accuracies that were pre-computed during the main grid search, and record the improvement of the best sampled configuration over the initial draw. Repeating this procedure 256 times and averaging across the four LLMs produces the curves in Fig.˜2.
E.2 Random-Forest Regression Analysis Setup
@@
Every completed grid-search run — defined by a specific model, task, and random seed—serves as one training example for a model-agnostic importance analysis. Each example is encoded by the triple $(\beta,\text{top-}p,T)$ and labeled with its accuracy. We fit a RandomForestRegressor from scikit-learn [44] with 100 trees that have unrestricted depth. Impurity-based Gini importances rank the hyperparameters as $\beta$ (0.41), top-p (0.32), and temperature (0.27), as shown in Fig.˜2.
@@
Table A5: Hyperparameter configuration by task. AIME and Count Down 4 use grid-search ranges; GPQA-Diamond and LiveCodeBench share a single universal setting.
@@
$\beta$ $\{\,\tfrac{1}{4},\tfrac{1}{2},1,2,4,8\}$ $\{\,\tfrac{1}{4},\tfrac{1}{2},1,2,4,8\}$ 11 11
@@
E.3 Setup for Case Study
@@
During inference, we prepend this blended vector to each of the 10-shot demonstrations and feed the embeddings directly to the model. For the single-prompt baseline we calculate the average accuracy of the prompts as the expectation of randomly choosing any single prompt, whereas the blended prompt is evaluated once; consequently, Table˜2 reflects an identical number of forward passes per model. Because all other factors remain fixed, any performance difference isolates the effect of the prompt representation itself.
Appendix F Hyperparameters
Table˜A5 lists the full search space for AIME and Count Down 4, along with the universal settings used for GPQA-Diamond and LiveCodeBench. All searches use five random seeds; reported results are seed-averaged.
Appendix G Implementation Details